OpenStudy (anonymous):

Need Help Solve 3|2x − 1| − 4 = 29 for x

OpenStudy (anonymous):

|x| = n can be expressed as 2 equations x=n and -x=n

OpenStudy (anonymous):

But how will that be the answer?

OpenStudy (anonymous):

it will help later in the problem

OpenStudy (anonymous):

3|2x − 1| − 4 = 29 for x 6x-3-4=29 6x-3=29+4 6x-3=33 6x=33+3 6x=36 x=36/6 x=6

OpenStudy (anonymous):

3[12-1]-4=29 36-3-4=29 33-4=29